# Motion Manifold Flow Primitives for Language-Guided Trajectory Generation

### v1

@@ Abstract
Keywords: LfD, movement primitives, manifold, flow-based models
@@ 1 Introduction
Past successes in text-based generative models for applications ranging from image synthesis [1, 2] and 3D scene construction [3, 4] to human motion generation [5, 6] can broadly be traced to two technical advancements: (i) large-scale pre-trained text embedding models based on transformer architectures [7, 8], and (ii) efficient training methods for probability density flow-based models that are capable of learning complex conditional density functions, e.g., diffusion [9, 10], score-based [11], and continuous normalizing flow models [12, 13, 14]. In this paper, we address the problem of generating a set of robot motions from a user-provided text instruction.
@@
Instead, leveraging a low-dimensional manifold structure can be an effective solution. Assuming the set of demonstration trajectories lies on a lower-dimensional manifold embedded in the trajectory space (the manifold hypothesis), recent works such as Motion Manifold Primitives (MMP) [15] and Task-Conditional Variational Autoencoders (TCVAE) [16] exploit this structure. These methods use an autoencoder framework, where a decoder transforms a low-dimensional latent space distribution (e.g., Gaussian or Gaussian Mixture Model) to the task-conditional data distribution, given a task parameter. These models can be used by regarding texts as task parameters (Figure 1 Middle). However, text-conditional trajectory distributions are inherently complex, with varying numbers of modalities and volumes of support based on the input text. Our studies indicate that a single decoder network struggles to learn these complex conditional distributions.
@@
In this paper we propose the Motion Manifold Flow Primitive (MMFP) framework, which combines the advantage of flow-based models capable of learning complex conditional distributions with motion manifold primitives capable of encoding and generating low-dimensional manifold of trajectories. Our proposed method consists of two steps. We first learn a decoder that parametrizes the motion manifold. In the low-dimensional latent coordinate space, we then train a flow-based model by using the recent flow matching algorithm of [12]. Specifically, we design proper regularization terms to ensure the smoothness of the generated trajectories and robustness against unstructured text variations that preserve meanings. As illustrated in Figure 1 (Lower), given a text, the latent flow-based model first transforms the low-dimensional latent Gaussian to the text-conditional latent space distribution, and then the decoder maps it to the trajectory space distribution.
@@
We evaluate the performance of MMFP through two text-based trajectory generation tasks: (i) a bottle’s SE(3) trajectory generation for pouring task and (ii) a 7-DoF robot arm’s waving trajectory generation task. Notably, with only 10 and 30 demonstration trajectories for each respective task, MMFP shows excellent performance. In contrast, diffusion and flow-based models trained directly in the high-dimensional trajectory space [9, 12, 13] and existing motion manifold primitives-based methods [16, 15] fail to generate successful trajectories.
@@ 2.1 Movement primitives
There are many studies on movement primitives, e.g., dynamic movement primitives [17, 18, 19, 20, 21, 22, 23, 24, 25, 26], stable dynamical systems [27, 28, 29, 30, 31, 32, 33, 34, 35], methods to represent diverse motions [36, 37, 38, 39, 40, 16, 15]. Of particular relevance to our problem is the task-conditional movement primitives capable of generating diverse trajectories that perform the given task. Text inputs can be transformed into text embedding vectors using pre-trained language models [7, 8], and these vectors can be treated as task parameters. Task-Parametrized Gaussian Mixture Model (TP-GMM) [36] is capable of generating new movements based on unseen task parameters. However, its emphasis is on a specific type of task parameters, such as frames of reference, making it unsuitable for handling text embedding vectors. Recent deep autoencoder-based methods such as Task-Conditional Variation Autoencoder (TCVAE) [16] and Motion Manifold Primitives (MMPs) [15] can take general types of task parameters. However, they show far less-than-desirable performance given the complexity of the text-conditional trajectory distribution.
@@ 2.2 Language-based generative models
Recent success of text-conditional deep generative models in image [1, 2, 41, 42, 43], human motion [5, 6, 44], and 3D scene generation tasks [3, 4] is not only attributed to transformer-based text embedding methods [7, 8, 45, 46, 47] and efficient training methods for flow-based conditional generative models [9, 10, 11, 12, 13, 14], but also relies heavily on the abundance of text-annotated training data. In contrast, text-based generation of robot arm motions has received relatively less attention, partly due to limited demonstration data. Notable exceptions include the open X-embodiment robot learning datasets [48] and generalist robotic transformers [49], which are trained using extensive text-annotated robot motion data. However, challenges persist in generating complex 6-DoF motions, confining their capabilities primarily to simpler actions like pick-and-place. Our objective sets us apart, as we aim to develop a language-based model that can generate complex 6-DoF motions for specific tasks, yet using a limited amount of demonstration data.
@@ 3.1 Autoencoder-based manifold learning
3.2 Continuous normalizing flow and flow matching
In this section, we introduce the continuous normalizing flow [57] and its efficient training method, the flow matching algorithm [12]. Let $\{x_{i}\in{\cal X}\}_{i=1}^{N}$ be a set of data points sampled from the underlying probability density $q(x)$. Consider a time-dependant vector field $v:[0,1]\times{\cal X}\to T{\cal X}$ that leads to a flow $\phi:[0,1]\times{\cal X}\to{\cal X}$ via the following ordinary differential equation: $\frac{d}{dt}\phi_{t}(x)=v_{t}(\phi_{t}(x))$, where $\phi_{0}(x)=x$. Given a prior density at $t=0$ denoted by $p_{0}$, the flow $\phi_{t}$ leads to a probability density path for $t\in[0,1]$ [12]: $p_{t}(x)=p_{0}(\phi_{t}^{-1}(x))\det\big{(}\frac{\partial\phi_{t}^{-1}}{\partial x}(x)\big{)}$. Our objective is to learn a neural network model of $v_{t}(x)$ so that the flow of $v_{t}(x)$ transforms a simple prior density $p_{0}$ (e.g., Gaussian) to the target data distribution $p_{1}\approx q$. Then we can sample new data points by solving the ODE from $t=0$ to $t=1$ with initial points sampled from $p_{0}$.
[…]
Our particular interest is a conditional density function $p(x|c)$ for a condition variable $c$ (e.g., text) which requires slight modifications. We model a neural network vector field $v_{t}(x,c)$ that takes an additional input $c$. Then, we minimize the following objective function:
$\mathbb{E}_{(x_{1},c)\sim q(x,c),\>t\sim{\cal U}[0,1],\>x\sim p_{t}(x|x_{1})}\big{[}\|v_{t}(x,c)-u_{t}(x|x_{1})\|^{2}\big{]},$ (1)
@@ 4 Language-based Robot Motion Generation
Figure 2: The procedure of motion generation in MMFP: (i) the Sentence-BERT encodes a free-form text into a vector $c$, (ii) the text embedding model $h_{\theta}$ maps $c$ to a text embedding vector $\tau$, (iii) we solve the ODE $\dot{z}=v_{t}(z,\tau;\theta)$ from $t=0$ to $t=1$ with an initial value $z_{0}$ sampled from Gaussian ${\cal N}(z|0,I)$ and obtain $z_{1}\in{\cal Z}$, and (iv) the motion decoder $f_{\theta}$ maps $z_{1}$ to a trajectory $x=(q_{1},\ldots,q_{T})$.
We begin this section with some notations. Let ${\cal Q}$ be a configuration space. Denote a sequence of configurations by $x=(q_{1},\ldots,q_{T})$, called a trajectory, for $q_{i}\in{\cal Q}$ and some positive integer $T>0$. The number of configurations $T$ is fixed throughout. We assume that the time interval between $q_{i}$ and $q_{i+1}$ is constant as $dt$, so the total time is $(T-1)\times dt$. We will denote the trajectory space by ${\cal X}={\cal Q}^{T}$. We use a pre-trained text encoder, the Sentence-BERT [58], without fine-tuning and it encodes free-form texts into 768-dimensional vectors; these encoded vectors are denoted by $c$. We assume that we are provided with a set of demonstration trajectories, each of which is annotated with $M$ different texts, i.e., the dataset is ${\cal D}=\{(x_{i},\{c_{ij}\}_{j=1}^{M})\}_{i=1}^{N}$ for $x_{i}\in{\cal X}$ and $c_{ij}\in\mathbb{R}^{768}$. We will use the same symbol $\theta$ to represent the learnable parameters of different neural networks with a slight abuse of notation, and denote a parametric neural network function by $f_{\theta}(x)$ or $f(x;\theta)$.
The overall framework of our method is shown in Figure 2. Our framework consists of two modules: (i) motion manifold model and (ii) latent space flow-based models. Below, we describe how to train each module.
Smooth motion manifold learning: The motion manifold model consists of two neural networks, an encoder $g_{\theta}:{\cal X}\to{\cal Z}$ and a decoder $f_{\theta}:{\cal Z}\to{\cal X}$, where ${\cal Z}=\mathbb{R}^{m}$ is the latent space. Given a suitable distance metric $d(\cdot,\cdot)$ in ${\cal X}$, we train an autoencoder as follows:
$\min_{\theta}\frac{1}{N}\sum_{i=1}^{N}d^{2}(x_{i},f_{\theta}(g_{\theta}(x_{i})))+\eta\|g_{\theta}(x_{i})\|^{2}+\delta{\cal E}(f_{\theta},g_{\theta}),$ (2)
@@
where $z=\alpha g_{\theta}(x_{i})+(1-\alpha)g_{\theta}(x_{j})$ and $\alpha\sim{\cal U}[-0.4,1.4]$ and $x_{i},x_{j}$ are sampled from the dataset. Then the trained decoder parametrizes an $m$-dimensional motion manifold embedded in the trajectory space (see section 3.1).

### v2

@@ Abstract
keywords:
Learning from Demonstration (LfD), movement primitives, manifold learning, autoencoders, flow-based models
@@ 1 Introduction
Past successes in text-based generative models for applications ranging from image synthesis (Rombach et al., 2022; Saharia et al., 2022) and 3D scene construction (Poole et al., 2022; Lin et al., 2023) to human motion generation (Zhang et al., 2022; Kim et al., 2023) can broadly be traced to two technical advancements: (i) large-scale pre-trained text embedding models based on transformer architectures (Devlin et al., 2018; Liu et al., 2019), and (ii) efficient training methods for probability density flow-based models that are capable of learning complex conditional density functions, e.g., diffusion (Ho et al., 2020; Song et al., 2020a), score-based (Song et al., 2020b), and continuous normalizing flow models (Lipman et al., 2022; Chen and Lipman, 2023; Tong et al., 2023). In this paper, we address the problem of generating a set of robot motions from a user-provided text instruction.
@@
Instead, leveraging a low-dimensional manifold structure can be an effective solution. Assuming the set of demonstration trajectories lies on a lower-dimensional manifold embedded in the trajectory space (the manifold hypothesis), recent works such as Motion Manifold Primitives (MMP) (Lee et al., 2023) and Task-Conditional Variational Autoencoders (TCVAE) (Noseworthy et al., 2020) exploit this structure. These methods use an autoencoder framework, where a decoder transforms a low-dimensional latent space distribution (e.g., Gaussian or Gaussian Mixture Model) to the task-conditional data distribution, given a task parameter. These models can be used by regarding texts as task parameters (Figure 1 Middle). However, text-conditional trajectory distributions are inherently complex, with varying numbers of modalities and volumes of support based on the input text. Our studies indicate that a single decoder network struggles to learn these complex conditional distributions.
@@
In this paper we propose the Motion Manifold Flow Primitive (MMFP) framework, which combines the advantage of flow-based models capable of learning complex conditional distributions with motion manifold primitives capable of encoding and generating low-dimensional manifold of trajectories. Our proposed method consists of two steps. We first learn a decoder that parametrizes the motion manifold. In the low-dimensional latent coordinate space, we then train a flow-based model by using the recent flow matching algorithm of (Lipman et al., 2022). Specifically, to improve our model’s robustness against unseen text variations, we introduce a regularization term using the Large Language Model ChatGPT, significantly enhancing its robustness. As illustrated in Figure 1 (Lower), given a text, the latent flow-based model first transforms the low-dimensional latent Gaussian to the text-conditional latent space distribution, and then the decoder maps it to the trajectory space distribution.
@@
We evaluate the performance of MMFP through two text-based trajectory generation tasks: (i) a bottle’s SE(3) trajectory generation for pouring task and (ii) a 7-DoF robot arm’s waving trajectory generation task. Notably, with only 10 and 30 demonstration trajectories for each respective task, MMFP shows excellent performance. In contrast, diffusion and flow-based models trained directly in the high-dimensional trajectory space (Ho et al., 2020; Lipman et al., 2022; Chen and Lipman, 2023) and existing motion manifold primitives-based methods (Noseworthy et al., 2020; Lee et al., 2023) fail to generate successful trajectories.
The main contributions of our work can be summarized as follows:
  • We propose MMFP that simultaneously addresses the challenges of small dataset, high-dimensionality, and complex text-dependency of trajectory distributions by integrating motion manifolds and flows, thereby achieving what neither approach can accomplish on its own;
  • We propose a regularization term that leverages the recent Large Language Model ChatGPT, significantly improving robustness against unseen and unstructured text variations.
@@ 2.1 Movement primitives
There are many studies on movement primitives, e.g., dynamic movement primitives (Saveriano et al., 2021; Ijspeert et al., 2013, 2001, 2002; Schaal et al., 2007; Pervez et al., 2017a; Fanger et al., 2016; Umlauft et al., 2017; Pervez et al., 2017b; Bahl et al., 2020), stable dynamical systems (Khansari-Zadeh and Billard, 2011; Neumann et al., 2013; Khansari-Zadeh and Billard, 2014; Lemme et al., 2014; Neumann and Steil, 2015; Blocher et al., 2017; Figueroa and Billard, 2018; Sindhwani et al., 2018; Kolter and Manek, 2019), methods to represent diverse motions (Calinon, 2016; Duque et al., 2019; Yang et al., 2018; Chernova and Veloso, 2007; Paraschos et al., 2013; Noseworthy et al., 2020; Lee et al., 2023; Lee, 2024). Of particular relevance to our problem is the task-conditional movement primitives. Text inputs can be transformed into text embedding vectors using pre-trained language models (Devlin et al., 2018; Liu et al., 2019), and these vectors can be treated as task parameters. Task-Parametrized Gaussian Mixture Model (TP-GMM) (Calinon, 2016) is capable of generating new movements based on unseen task parameters. However, its emphasis is on a specific type of task parameters, such as frames of reference, making it unsuitable for handling text embedding vectors.
Recent deep learning methods that adopt the conditional autoencoder architectures (Noseworthy et al., 2020; Hristov and Ramamoorthy, 2021; Lee et al., 2023) can take general types of task parameters and encode and generate diverse trajectories. However, as demonstrated in our later experiments, they show poor performance given the complexity of the text-conditional trajectory distributions.
@@ 2.2 Language-based generative models
Recent success of text-conditional deep generative models in image (Rombach et al., 2022; Saharia et al., 2022; Ruiz et al., 2023; Zhang et al., 2023; Balaji et al., 2022), human motion (Zhang et al., 2022; Kim et al., 2023; Tevet et al., 2022), and 3D scene generation tasks (Poole et al., 2022; Lin et al., 2023) is not only attributed to transformer-based text embedding methods (Devlin et al., 2018; Liu et al., 2019; Wang et al., 2022; Li et al., 2023; Beltagy et al., 2020) and efficient training methods for flow-based conditional generative models (Ho et al., 2020; Song et al., 2020a, b; Lipman et al., 2022; Chen and Lipman, 2023; Tong et al., 2023), but also relies heavily on the abundance of text-annotated training data.
In contrast, text-based generation of robot arm motions has received relatively less attention, partly due to limited demonstration data. Notable exceptions include the open X-embodiment robot learning datasets (Collaboration, 2023) and generalist robotic transformers (Brohan et al., 2022), which are trained using extensive text-annotated robot motion data. However, challenges persist in generating complex 6-DoF motions, confining their capabilities primarily to simpler actions like pick-and-place. Our objective sets us apart, as we aim to develop a language-based model that can generate complex 6-DoF motions for specific tasks, yet using a limited amount of demonstration data.
@@ 3.1 Autoencoder-based manifold learning
In this section, we briefly introduce an autoencoder and its manifold learning perspective (Arvanitidis et al., 2017; Lee et al., 2021, 2022a, 2022b; Jang et al., ; Lee, 2023; Lee and Park, 2023; Lim et al., 2024). Consider a high-dimensional data space ${\cal X}$ and a set of data points ${\cal D}=\{x_{i}\in{\cal X}\}_{i=1}^{N}$. We adopt the manifold hypothesis that the data points $\{x_{i}\}$ lie approximately on some lower-dimensional manifold ${\cal M}$ in ${\cal X}$. Suppose ${\cal M}$ is an $m$-dimensional manifold and let ${\cal Z}$ be a latent space $\mathbb{R}^{m}$. An encoder is a mapping $g:{\cal X}\to{\cal Z}$ and a decoder is a mapping $f:{\cal Z}\to{\cal X}$. These are often approximated with deep neural networks and trained to minimize the reconstruction loss
$\frac{1}{N}\sum_{i=1}^{N}d^{2}(f\circ g(x_{i}),x_{i})$ (1)
given a distance metric $d(\cdot,\cdot)$ on ${\cal X}$. We note that, given a sufficiently low reconstruction error, all the data points $\{x_{i}\}$ should lie on the image of the decoder $f$. Under some mild conditions – that are (i) $m$ is lower than the dimension of the data space ${\cal X}$ and (ii) $f$ is smooth and its Jacobian $\frac{\partial f}{\partial z}(z)\in\mathbb{R}^{dim({\cal X})\times m}$ is full rank everywhere –, the image of $f$ is an $m$-dimensional differentiable manifold embedded in ${\cal X}$, i.e., the decoder produces a lower-dimensional manifold where the data points approximately lie.
In this paper, we will treat the trajectory space as ${\cal X}$ – where each element corresponds to a discrete-time trajectory – and will attempt to learn a lower-dimensional manifold of these trajectories to address the challenges of small dataset size and high dimensionality.
3.2 Continuous normalizing flow and flow matching
In this section, we introduce the continuous normalizing flow (Chen et al., 2018) capable of fitting complex probability distributions and its efficient training method, the flow matching algorithm (Lipman et al., 2022). Let $\{x_{i}\in{\cal X}\}_{i=1}^{N}$ be a set of data points sampled from the underlying probability density $q(x)$. Consider a non-autonomous vector field $v:[0,1]\times{\cal X}\to T{\cal X}$ that leads to a flow $\phi:[0,1]\times{\cal X}\to{\cal X}$ via the following ordinary differential equation:
$\frac{d}{ds}\phi_{s}(x)=v_{s}(\phi_{s}(x)),$ (2)
where $\phi_{0}(x)=x$. Given a prior density at $s=0$ denoted by $p_{0}$, the flow $\phi_{s}$ leads to a probability density path for $s\in[0,1]$ (Lipman et al., 2022):
$p_{s}(x)=p_{0}(\phi_{s}^{-1}(x))\det\big{(}\frac{\partial\phi_{s}^{-1}}{\partial x}(x)\big{)}.$ (3)
Our objective is to learn a neural network model of $v_{s}(x)$ so that the flow of $v_{s}(x)$ transforms a simple prior density $p_{0}$ (e.g., Gaussian) to the target data distribution $p_{1}\approx q$. Then we can sample new data points by solving the ODE, $x^{\prime}=v_{s}(x)$, from $s=0$ to $s=1$ with initial points sampled from $p_{0}$.
The standard maximum log-likelihood training requires expensive numerical ODE simulations (Chen et al., 2018), instead we introduce an efficient simulation-free approach for training $v_{s}(x)$ (Lipman et al., 2022). In the first step, we design a conditional probability path $p_{s}(x|x_{1})$ for $x_{1}\sim q(x)$ such that $p_{0}(x|x_{1})=p_{0}(x)$ and $p_{1}(x|x_{1})$ is concentrated around $x=x_{1}$. For example, Lipman et al. (2022) suggests an Optimal Transport (OT) Gaussian probability path $p_{s}(x|x_{1})={\cal N}(x|sx_{1},(1-(1-\sigma_{min})s)^{2}I)$ where $\sigma_{min}$ is set to be small. Theorem 3 in (Lipman et al., 2022) states that $u_{s}(x|x_{1})=\frac{x_{1}-(1-\sigma_{min})x}{1-(1-\sigma_{min})s}$ generates the Gaussian path $p_{s}(x|x_{1})$. Then, Lipman et al. (2022) shows that minimizing the following flow matching objective function
$\mathbb{E}_{x_{1}\sim q(x),\>s\sim{\cal U}[0,1],\>x\sim p_{s}(x|x_{1})}\big{[}\|v_{s}(x)-u_{s}(x|x_{1})\|^{2}\big{]},$ (4)
@@ 3.2 Continuous normalizing flow and flow matching
Our particular interest is a conditional density function $p(x|c)$ for a condition variable $c$ (e.g., text) which requires slight modifications. We model a neural network vector field $v_{s}(x,c)$ that takes an additional input $c$. Then, we minimize the following objective function:
$\mathbb{E}_{(x_{1},c)\sim q(x,c),\>s\sim{\cal U}[0,1],\>x\sim p_{s}(x|x_{1})}\big{[}\|v_{s}(x,c)-u_{s}(x|x_{1})\|^{2}\big{]},$ (5)
@@
We note that this model can represent the complex dependency of the distribution on the condition variable, e.g., when the number of modalities or the support volumes of $p(x|c)$ vary significantly as $c$ changes. This is because, while a single continuous neural network model is not capable of representing $p(x|c)$ when it changes dramatically depending on $c$, a vector field $v_{s}(x,c)$, with an additional input $s$, can continuously transform the distribution as $s$ progresses from $s=0$ to $s=1$, thereby representing complex $p(x|c)$. In this work, the inherent complex text-dependency of the motion distribution will be captured by leveraging these flow-based models.
@@ 4 Language-based Robot Motion Generation
Figure 2: The procedure of motion generation in MMFP: (i) the Sentence-BERT encodes a free-form text into a vector $c$, (ii) the text embedding model $h_{\theta}$ maps $c$ to a text embedding vector $\tau$, (iii) we solve the ODE $z^{\prime}=v_{s}(z,\tau;\theta)$ from $s=0$ to $s=1$ with an initial value $z_{0}$ sampled from Gaussian ${\cal N}(z|0,I)$ and obtain $z_{1}\in{\cal Z}$, and (iv) the motion decoder $f_{\theta}$ maps $z_{1}$ to a trajectory $x=(q_{1},\ldots,q_{T})$.
We begin this section with some notations. Let ${\cal Q}$ be a configuration space. Denote a sequence of configurations by $x=(q_{1},\ldots,q_{T})$, called a trajectory, for $q_{i}\in{\cal Q}$ and some positive integer $T>0$. The number of configurations $T$ is fixed throughout. We assume that the time interval between $q_{i}$ and $q_{i+1}$ is constant as $dt$, so the total time is $(T-1)\times dt$. We will denote the trajectory space by ${\cal X}={\cal Q}^{T}$. We use a pre-trained text encoder, the Sentence-BERT (Reimers and Gurevych, 2019), without fine-tuning and it encodes free-form texts into 768-dimensional vectors; these encoded vectors are denoted by $c$. We assume that we are provided with a set of demonstration trajectories, each of which is annotated with $M$ different texts, i.e., the dataset is ${\cal D}=\{(x_{i},\{c_{ij}\}_{j=1}^{M})\}_{i=1}^{N}$ for $x_{i}\in{\cal X}$ and $c_{ij}\in\mathbb{R}^{768}$. We will use the same symbol $\theta$ to represent the learnable parameters of different neural networks with a slight abuse of notation, and denote a parametric neural network function by $f_{\theta}(x)$ or $f(x;\theta)$.
The overall framework of our method is shown in Figure 2. Our framework consists of two modules: (i) the motion manifold, which addresses the challenges of small dataset size and the high dimensionality of the trajectory space ${\cal Q}^{T}$, and (ii) latent space flows, which are capable of capturing the complex text-dependency of the motion distribution. In the subsequence sections, we describe the training process for each module.
4.1 Smooth Motion Manifold Learning
The motion manifold model consists of two neural networks, an encoder $g_{\theta}:{\cal X}\to{\cal Z}$ and a decoder $f_{\theta}:{\cal Z}\to{\cal X}$, where ${\cal Z}=\mathbb{R}^{m}$ is the latent space. Given a suitable distance metric $d(\cdot,\cdot)$ in ${\cal X}$, we train an autoencoder as follows:
$\small\min_{\theta}\frac{1}{N}\sum_{i=1}^{N}d^{2}(x_{i},f_{\theta}(g_{\theta}(x_{i})))+\eta\|g_{\theta}(x_{i})\|^{2}+\delta{\cal E}(f_{\theta},g_{\theta}),$ (6)
@@ 4.1 Smooth Motion Manifold Learning
where $f_{\theta}(z)=(f^{1}_{\theta}(z),\ldots,f^{T}_{\theta}(z))\in{\cal Q}^{T}$. The latent value $z$ in this expectation is sampled with augmentation as $z=\alpha g_{\theta}(x_{i})+(1-\alpha)g_{\theta}(x_{j})$, where $\alpha\sim{\cal U}[-0.4,1.4]$ and $x_{i},x_{j}$ are sampled from the dataset, to extend the regularization effect to regions where data is not available, adopting (Lee et al., 2022b). We have empirically observed that with sufficiently low-dimensional latent spaces and smooth activation functions, the decoder $f_{\theta}$ satisfies the injective immersion condition, and the trained decoder parametrizes an $m$-dimensional motion manifold embedded in the trajectory space (see the preliminary section on manifold learning).
4.2 Robust Text-conditional Latent Flow Learning